OpenStudy (anonymous):

** PLEASE HELP ** Write the equation of the line that is perpendicular to the line y = 3/2x + 4 and passes through the point (-6, 3). A. y = 3/2x - 1 B. y = 3/2x + 3 C. y = -2/3x - 1 D. y = -2/3x + 3

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (anonymous):

Im completely lost for this one

jimthompson5910 (jim_thompson5910):

The slope of y = 3/2x + 4 is ????

OpenStudy (anonymous):

0 ?

jimthompson5910 (jim_thompson5910):

nope

jimthompson5910 (jim_thompson5910):

hint: y = mx+b has a slope of m

OpenStudy (anonymous):

3/2

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

flip the fraction and flip the sign to get ______

OpenStudy (anonymous):

what... ,_,

jimthompson5910 (jim_thompson5910):

you have the fraction 3/2 right?

jimthompson5910 (jim_thompson5910):

to find the perpendicular slope, you flip that fraction and you flip the sign

OpenStudy (anonymous):

-2/3 ?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now solve for b \[\Large y = mx + b\] \[\Large y = -\frac{2}{3}x + b\] \[\Large 3 = -\frac{2}{3}(-6) + b\] \[\Large b = ???\]

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

m = -2/3 b = -1 the final answer is \[\Large y = -\frac{2}{3}x - 1\]

OpenStudy (anonymous):

You're such a big help, thank you!!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Would you mind helping me with one more ?

jimthompson5910 (jim_thompson5910):

sure I can help

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -1/3x - 6 y = -1/3x - 14 y = 3x - 6 y = 3x - 14

jimthompson5910 (jim_thompson5910):

First I would solve for y to get it into y = mx+b form (slope intercept form)

jimthompson5910 (jim_thompson5910):

3x - y = -3 -y = -3-3x -y = -3x-3 y = (-3x-3)/(-1) y = 3x + 3 with me so far?

OpenStudy (anonymous):

Im with you (:

jimthompson5910 (jim_thompson5910):

The slope of y = 3x + 3 is m = 3 the slope of ANY parallel line to this given one will also have a slope of 3 any two parallel lines have equal slopes

jimthompson5910 (jim_thompson5910):

y = mx + b y = 3x + b -2 = 3*4 + b b = ???

OpenStudy (anonymous):

-14

jimthompson5910 (jim_thompson5910):

so your answer is y = 3x - 14

OpenStudy (anonymous):

You're the best!